OpenStudy (anonymous):

Solve triangle PQR, given:, q = 38.2, and r = 39.6

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

do u have anything else? or is this question complete?

OpenStudy (anonymous):

u need a third known element too, i-e a side or an angle

OpenStudy (anonymous):

the angle is Q=71 degrees

OpenStudy (anonymous):

\[\frac{ q }{ \sin(Q) }=\frac{ p }{ \sin(P) }=\frac{ r }{ \sin(R) }\] \[\frac{ q }{ \sin(Q) }=\frac{ r }{ \sin(R) }\] \[\frac{ 38.2 }{ \sin(71) }=\frac{ 39.6}{ \sin(R) }\] solve for R

OpenStudy (anonymous):

so do I cross multiply?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

sorry how do you cross multiply with this case?

OpenStudy (anonymous):